If the pH at the half equivalence point is 4.2, what is the pKa of the acid? Half equivalence point is also called the midpoint in the text.
Answer #1

At half equivalence point, concentration of acid is equal to concentration of conjugate base

So,

[A-] = [HA]

Using:

pH = pKa + log [A-]/[HA]

4.2 = pKa + log (1)

4.2 = pKa + 0

pKa = 4.2

Answer: 4.2